A 0.30 kg block is pulled up a 25 degree incline by an ideal string under tension; the string is draped over an ideal pulley. Tension in the string is created by a second block of equal mass, suspended by the section of the string directly below the pulley. The coefficient of friction between the first block and the incline is 0.40. After a slight push, what would be (a) the acceleration of the blocks, and (b) the tension in the string?
Answer #1

here,

the mass of block , m = 0.3 kg

theta = 25 degree

the coefficient of friction , uk = 0.4

a)

the acceleration of the blocks , a = (m * g - m * g * sin(theta) - uk * m * g * cos(theta)) /(m + m)

a = 9.81 * ( 1 - sin(25) - 0.4 * cos(25)) /2 m/s^2

a = 1.06 m/s^2

b)

for hanging block

m * g - T = m * a

T = 0.3 * ( 9.81 - 1.06) N

T = 2.63 N

the tension in the string is 2.63 N